A collection of 40 coins consisting of dimes and quarters has a value of $5.95. Find the number of each kind of coin.
Answer by mananth(16946)

dimes ----x
quarters --y
x+y=40
10x+25y=595
x numbers dimes
y numbers quarters

x +y=40 .............1
10 x + 25 y = 595 .............2
Eliminate y
multiply (1)by -25
Multiply (2) by 1
-25 x -25 y = -1000
10 x + 25 y = 595
Add the two equations
-15 x = -405.00
/ -15
x = 27
plug value of x in (1)
1 x + y = 40
27 + y = 40
y = 40
y = 13
y = 13
dimes 27
quarters 13
m.ananth@hotmail.ca